8^C4/12^C4

8C4 over 12C4=
A combination is a selection of objects from a group of objects without regard to order and is denoted by nCr.
I've tried dividing it out and that is not the response for this statistics question.

Answer by Alan3354(69443)

8^C4/12^C4

8C4 over 12C4 = 8C4/12C4
----------
The ^ implies exponentiation.
8C4 = 8!/(4!*(8-4)!) = 8*7*6*5/24 = 70
12C4 = 12!/((12-4)!*4!) = 495
--------------
70/495 = 14/99